### v1

@@ Appendix A Autoregressive Poisson Stein-like lemma
Proof of driven autoregressive Poisson Stein-like lemma 1.
Let $\boldsymbol{\mathsf{Y}}\in\mathbb{R}_{+}^{T}$ random observations under the driven generalized autoregressive model (7) with ground truth reproduction number $\overline{\boldsymbol{\mathsf{X}}}\in\mathbb{R}_{+}^{T}$ and memory functions $\{\Psi_{t},\,t=1,\ldots,T\}$ satisfying Assumption 2, following a scaled Poisson distribution (11) of time-varying modulus $\boldsymbol{\alpha}\in(\mathbb{R}_{+}^{*})^{T}$, and a function $\Theta:\mathbb{R}_{+}^{T}\rightarrow\mathbb{R}$ satisfying Assumption 1. First, by Assumption 1, the expectation in the left-hand side of Equation (32) is well defined. Hence, using the discrete probabilistic density function of the scaled Poisson distribution leads to
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ (53)
$=\sum_{\mathsf{k}_{1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})},$
@@
$\sum_{\mathsf{k}_{t}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!$ $\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}$
$=$ $\,\prod_{s=1}^{t-1}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}$ (54)
×$\times\sum_{k_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}$
where $\mathsf{G}(\boldsymbol{\mathsf{Y}})=$
$\!\!\!\!\!\sum_{\mathsf{k}_{t+1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\prod_{u=t+1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}.$ (55)
@@
$\begin{split}\sum_{\mathsf{k}_{t}=0}^{\infty}&\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\\ =&\sum_{\mathsf{k}_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\times(\mathsf{k}_{t}+1)\times\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{{\mathsf{k}_{t}+1}}}{(\mathsf{k}_{t}+1)!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\\ \end{split}$ (56)
@@
$\begin{split}\sum_{\mathsf{k}_{t}=0}^{\infty}&\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\\ &=\sum_{\mathsf{k}_{t}=0}^{\infty}\mathsf{G}^{-t}(\boldsymbol{\mathsf{Y}})\mathsf{k}_{t}\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{{\mathsf{k}_{t}}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\end{split}$ (57)
@@
The above discrete counterpart of integration by part on variable $\mathsf{k}_{t}$ performed in Equations (56) and (57) can be seen as an application of the standard Poisson Stein’s lemma counterpart [41, 38] on the function $\mathsf{G}$. Though, this reformulation is not enough to design Poisson Unbiased Risk Estimates for the driven autoregressive model (7) as, contrary to the case of standard Poisson Unbiased Risk Estimate, neither the function $\mathsf{G}$ nor its translate versions $\mathsf{G}^{-t}$ can be reformulated easily as expectations of tractable quantities.
@@
where the change of variable $\boldsymbol{\mathsf{Y}}\rightarrow\widetilde{\boldsymbol{\mathsf{Y}}}$ only affects the term in $\Theta$ but not the Poisson densities. Finally, injecting (59) into (53), yields $\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\simeq$
$\sum_{\mathsf{k}_{1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta^{-t}(\boldsymbol{\mathsf{Y}})\mathsf{k}_{t}\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}$ (60)
@@ Appendix B Autoregressive Poisson Unbiased Risk Estimate
$\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})-\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}=\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}$
$-2\left\langle\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}),\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rangle+\left\lVert\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}.$ (61)
The first term in the right-hand side of (61) only depends on observations and hyperparameters, and hence appears as is in $\mathsf{APURE}^{\mathcal{P}}$ in Equation (LABEL:eq:apure). The second and third terms depend on the inaccessible ground truth $\overline{\boldsymbol{\mathsf{X}}}$, and hence need to be reformulated, taking care of not introducing any bias. Considering the second term,
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}$ $\left[\left\langle\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}),\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rangle\right]$
$=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right].$
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}&\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\\ &\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\right)^{-t}\mathsf{Y}_{t}\right]\end{split}$ (62)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}&\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\\ &\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]\end{split}$ (63)
in which the reader recognizes the expression of the second term in the definition of $\mathsf{APURE}^{\mathcal{P}}$ in Equation (LABEL:eq:apure). As for the third term of Equation (61), it writes
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\overline{\mathsf{X}}_{t}\right]=\,\,\,&\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\\ \underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}&\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\\ \underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}&\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]\end{split}$ (68)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left\lVert\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2}\right]$ $=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\overline{\mathsf{X}}_{t}^{2}\right]$ (69)
$=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right].$ (70)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ =$\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]$ (71)
=$\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ (72)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]&\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\\ &\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{(\mathsf{Y}_{t}-\alpha_{t})\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right].\end{split}$ (73)

### v3

@@ Appendix A Autoregressive Poisson Stein-like lemma
Proof of generalized time-varying autoregressive Poisson Stein-like lemma 1.
Let $\boldsymbol{\mathsf{Y}}\in\mathbb{R}_{+}^{T}$ random observations under the generalized time-varying autoregressive model (7) with ground truth reproduction number $\overline{\boldsymbol{\mathsf{X}}}\in\mathbb{R}_{+}^{T}$ and memory functions $\{\Psi_{t},\,t=1,\ldots,T\}$ satisfying Assumption 2, following a scaled Poisson distribution (11) of time-varying modulus $\boldsymbol{\alpha}\in(\mathbb{R}_{+}^{*})^{T}$, and a function $\Theta:\mathbb{R}_{+}^{T}\rightarrow\mathbb{R}$ satisfying Assumption 1. First, by Assumption 1, the expectation in the left-hand side of Equation (30) is well defined. Hence, using the discrete probabilistic density function of the scaled Poisson distribution leads to
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]=\sum_{\mathsf{k}_{1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})},$ (50)
@@
$\begin{split}\sum_{\mathsf{k}_{t}=0}^{\infty}\!\ldots\!\!&\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}\\ &=\,\prod_{s=1}^{t-1}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}\times\sum_{k_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}\end{split}$ (51)
$\text{where}\quad\mathsf{G}(\boldsymbol{\mathsf{Y}})=\sum_{\mathsf{k}_{t+1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta(\boldsymbol{\mathsf{Y}})\prod_{u=t+1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}.$ (52)
@@
$\sum_{\mathsf{k}_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}=\sum_{\mathsf{k}_{t}=0}^{\infty}\mathsf{G}(\boldsymbol{\mathsf{Y}})\times(\mathsf{k}_{t}+1)\times\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{{\mathsf{k}_{t}+1}}}{(\mathsf{k}_{t}+1)!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}$ (53)
@@
$\sum_{\mathsf{k}_{t}=0}^{\infty}$ $\mathsf{G}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{t}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}=\sum_{\mathsf{k}_{t}=0}^{\infty}\mathsf{G}^{-t}(\boldsymbol{\mathsf{Y}})\mathsf{k}_{t}\frac{(\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}}))^{{\mathsf{k}_{t}}}}{\mathsf{k}_{t}!}\mathrm{e}^{-\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})}$ (54)
@@
The above discrete counterpart of integration by part on variable $\mathsf{k}_{t}$ performed in Equations (53) and (54) can be seen as an application of the standard Poisson Stein’s lemma counterpart [50, 47] on the function $\mathsf{G}$. Though, this reformulation is not enough to design Poisson Unbiased Risk Estimates for the generalized time-varying autoregressive model (7) as, contrary to the case of standard Poisson Unbiased Risk Estimate, neither the function $\mathsf{G}$ nor its translate versions $\mathsf{G}^{-t}$ can be reformulated easily as expectations of tractable quantities.
@@
where the change of variable $\boldsymbol{\mathsf{Y}}\rightarrow\widetilde{\boldsymbol{\mathsf{Y}}}$ only affects the term in $\Theta$ but not the Poisson densities. Finally, injecting (56) into (50), yields
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\Theta(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\simeq\sum_{\mathsf{k}_{1}=0}^{\infty}\!\ldots\!\!\sum_{\mathsf{k}_{T}=0}^{\infty}\!\!\Theta^{-t}(\boldsymbol{\mathsf{Y}})\mathsf{k}_{t}\prod_{s=1}^{T}\frac{(\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}}))^{\mathsf{k}_{s}}}{\mathsf{k}_{s}!}\mathrm{e}^{-\overline{\mathsf{X}}_{s}\Psi_{s}(\boldsymbol{\mathsf{Y}})}$ (57)
@@ Appendix B Autoregressive Poisson Unbiased Risk Estimate
$\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})-\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}=\left\lVert\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}-2\left\langle\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}),\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rangle+\left\lVert\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rVert^{2}_{2}.$ (58)
The first term in the right-hand side of (58) only depends on observations and hyperparameters, and hence appears as is in $\mathsf{APURE}^{\mathcal{P}}$ in Equation (31). The second and third terms depend on the inaccessible ground truth $\overline{\boldsymbol{\mathsf{X}}}$, and hence need to be reformulated, taking care of not introducing any bias. Considering the second term,
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left\langle\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}}),\overline{\boldsymbol{\mathsf{X}}}\odot\boldsymbol{\Psi}(\boldsymbol{\mathsf{Y}})\right\rangle\right]=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right].$
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\right)^{-t}\mathsf{Y}_{t}\right]\end{split}$ (59)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\Psi_{t}(\boldsymbol{\mathsf{Y}})\mathsf{Y}_{t}\right]\end{split}$ (60)
in which the reader recognizes the expression of the second term in the definition of $\mathsf{APURE}^{\mathcal{P}}$ in Equation (31). As for the third term of Equation (58), it writes
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})\overline{\mathsf{X}}_{t}\right]=\,\,\,\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\widehat{\mathsf{X}}_{t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\widehat{\mathsf{X}}_{t}^{-t}(\boldsymbol{\mathsf{Y}};\boldsymbol{\lambda})}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]$ (65)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left\lVert\overline{\boldsymbol{\mathsf{X}}}\right\rVert^{2}_{2}\right]=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\overline{\mathsf{X}}_{t}^{2}\right]=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right].$ (66)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ (67)
@@
$\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{\overline{\mathsf{X}}_{t}}{\Psi_{t}(\boldsymbol{\mathsf{Y}})}\overline{\mathsf{X}}_{t}\Psi_{t}(\boldsymbol{\mathsf{Y}})\right]$ =$\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\left(\frac{\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right)^{-t}\mathsf{Y}_{t}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathbb{E}_{\boldsymbol{\mathsf{Y}}}\left[\frac{(\mathsf{Y}_{t}-\alpha_{t})\mathsf{Y}_{t}}{\Psi_{t}^{2}(\boldsymbol{\mathsf{Y}})}\right].$ (68)
Equations (65) and (68), combined with the expansion provided in Equation (64) demonstrates the asymptotic unbiasedness of $\mathsf{APURE}^{\mathcal{P}}$ stated in Equation (34). ∎